8. Calculate the effective rate of interest on a certificate of deposit that offers a rate of interest of 7% per year, compouplease correct
Answer #1

Effective interest rate = (1 + i/n)^n - 1

= (1 + 7%/2)^2 - 1

= (1.035)^2 - 1

= 1.071225 - 1

= 0.071225

= 7.1225%

where,

i = nominal annual interest rate

n = number of compounding periods